If 1,w,w^2. are the cube roots of unity then the value of (1+w+w^2)(1+w^2+w)^6

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
x^3=1 has roots 1, w, w^2.
or x^3 – 1= 0
or (x – 1)(x^2 + x + 1)= 0
or x^2 + x + 1= 0 has roots  w, w^2.
so w^2 + w + 1= 0
hence (1+w+w^2)(1+w^2+w)^6= 0
kindly approve :)